OpenStudy (anonymous):

12 = (m2 + 5m)2 + (m2 + 5m)

OpenStudy (anonymous):

@Jared25 m = 5.7

OpenStudy (anonymous):

how did you find it?

OpenStudy (anonymous):

@jared25 12 = (m^2 + 5m) (2) + m^2 + 5m 12 = 3m^2 + 15m 12 - (3m^2 + 15m) = 3m^2 + 15m - (3m^2 + 15m) -3m^2 - 15m + 12 = 0 m = -b + sqrt (b^2 - 4ac)/2a m = -(- 15 + sqrt ((-15)^2 - 4 (-3)(12))/2(-3) m = 15 + sqrt (369)/-6 m = 5.7

OpenStudy (anonymous):

really sorry copied problem wrong its 12=(m^2+5m)^2 + m^2+5m